# Speed Limit Reduction Enhances Urban Worker Safety: Evidence from a Decade of Traffic Incidents in Santiago, Chile

### v1

@@ Abstract
Work-related traffic incidents significantly impact urban mobility and productivity. This study analyzes a decade of work-related traffic incident data (2012–2021) in Santiago, Chile, using records from a major social insurance company. We explore temporal, spatial, and demographic patterns in these incidents in urban and rural areas. We also evaluate the impact of a 2018 urban speed limit reduction law on incident injury severity. Using negative binomial regression, we assess how various factors, including the speed limit change, affect injury severity measured by prescribed medical leave days. Our analysis reveals distinct incident occurrence and severity patterns across different times, locations, and demographic groups. We find that motorcycles and cycles are associated with more severe injuries, with marginal effects of 26.94 and 13.06 additional days of medical leave, respectively, compared to motorized vehicles. Female workers tend to have less severe injuries, with an average of 7.57 fewer days of medical leave. Age is also a significant factor, with each year associated with 0.57 additional days of leave. Notably, the urban speed limit reduction is associated with a decrease of 4.26 days in prescribed medical leave for incidents in urban areas, suggesting that lower speed limits contribute to reduced injury severity in work-related traffic incidents. Our results provide insights for urban planning, transportation policy, and workplace safety initiatives, highlighting the potential benefits of speed management in urban areas for improving road safety and minimizing the economic impact of work-related incidents.
1 Introduction
Work trips are a crucial aspect of daily urban mobility, shaping the rhythm of cities and defining peak travel periods. The safety of these trips is important, as work-related crashes, collisions or falls impact individual health, urban productivity, and sustainability (United Nations, 2023). These incidents significantly impact urban dynamics by affecting workers, disrupting traffic, straining emergency services, and causing economic losses through reduced productivity (Chantith et al., 2021). In this paper, we use the term ’incident’ to refer to traffic incidents that may result in injury or property damage, avoiding the term ’accident’ which can imply unavoidability.
While disruptive to urban systems, traffic incidents are often underrepresented in transportation analyses. Traditional travel surveys rarely capture traffic incident data, and new data sources like mobile phone records, although helpful in understanding general mobility patterns (Widhalm et al., 2015; Graells-Garrido et al., 2023a), cannot identify specific traffic incidents due to anonymization and lack of meta-data. However, work-related incidents offer study opportunities, as they are more likely to be centrally recorded due to legal and insurance requirements.
@@ 1 Introduction
Various institutions prioritize worker safety and insurance in Santiago, overseen by the Superintendencia de Seguridad Social (SUSESO). This paper analyzes a dataset from one of these entities, Mutual de Seguridad de la Cámara Chilena de la Construcción (MUSEG), which covers approximately 31% of all insured workers in the country as of 2022. Our study examines around 33,000 traffic incidents in the Santiago metropolitan area and surrounding rural regions from 2012 to 2021, revealing temporal, geographical, and demographic patterns.
A significant event within our study period is the implementation of a new speed limit regulation on August 4th, 2018 (Law 21,103), which reduced the urban speed limit from 60 km/h to 50 km/h (Gobierno de Chile, 2019). While this law primarily affects cars, its impact extends to all transportation modes due to their interaction with vehicles on shared roads. We use this policy change as a natural experiment, employing a Negative Binomial regression to assess its effect on incident severity.
@@
    What are the temporal, spatial, and demographic patterns of work-related traffic incidents in Santiago?
@@
    How did the 2018 speed limit reduction impact the severity of work-related traffic incidents?
@@
Our analysis shows that after the law’s implementation, the severity of all traffic incident injuries decreased. Specifically, incidents required an average of 3.88 fewer days of medical leave. This finding suggests that the law’s impact goes beyond reducing incident frequency, allowing workers to return to their families and work activities sooner.
This study contributes to research on urban mobility safety. Less than 10% of road safety research focuses on low- and middle-income countries, despite these areas accounting for most road traffic deaths and injuries (Haghani et al., 2022). While previous studies have shown the effectiveness of speed limit reductions in various contexts (Elvik, 2012; Cleland et al., 2020; Milton et al., 2021), our work provides specific insights into the impact on work-related incidents in a Latin American urban context, as well as the factors related to the severity of such incidents.
Our findings align with emerging urban planning concepts such as “slow streets” (Bajada et al., 2023) and “15-minute cities” (Graells-Garrido et al., 2021), which promote safer, more accessible urban environments. Recent studies support these approaches: vehicle speeds decrease in shared spaces when both pedestrian and vehicle flows are high (Tzouras et al., 2023), while slow zones in Paris increased human activity by 44% (Salazar-Miranda et al., 2022). Our research extends these findings to work-related mobility, showing that lower urban speed limits contribute to reduced injury severity in all types of traffic incidents. These results collectively suggest that cities designed for slower-paced mobility not only enhance safety but also promote social interaction and urban vitality, offering promising avenues for future urban planning and policy development.
@@ 2 Related Work
Our study builds upon research in urban transportation safety, focusing on work-related incidents and the impact of speed limit regulations. This section reviews key literature in these areas.
@@ 2.2 Spatial Analysis of Traffic incidents
In the Latin American context, Mangones et al. (2024) compared crash risk on arterial roads, Bus Rapid Transit (BRT), and Buses with High Level of Service (BHLS) corridors in Bogotá, Colombia. Their findings suggest that different public transport infrastructures have varying impacts on safety. BRT networks provide lower crash rates for less severe collisions but increase injuries and fatalities. This underscores the importance of considering different road types and transport modes in safety analyses. Sukhai et al. (2021), studying South Africa, another Global South country, found that contextual effects create unequal road safety risks, with higher vulnerabilities for pedestrians, females, and public transport users linked to rurality and socio-economic factors.
@@ 2.3 Impact of Speed Limit Laws
The relationship between speed and road safety has been extensively studied. Elvik (2012) conducted a systematic review and meta-analysis, concluding that speeding is the most critical factor contributing to traffic injuries. Their work provides a model correlating speed enforcement with both the likelihood of crashes and the resulting health consequences.
Cleland et al. (2020) and Milton et al. (2021) examined the effects of 20 mph (32 km/h) zones and limits in various cities. Both studies found associations between lower speed limits and crash frequency and harm to road users. However, they also noted the need for more evidence on the impact of air pollution on livability.
@@ 2.4 Gender and Age Dynamics in Traffic incidents
González-Sánchez et al. (2021) investigated the influence of road type on traffic injury risk, considering gender and age. They found that men consistently have a higher risk of severe and fatal injuries across all modes of transport and road types. Havet et al. (2021) explored how determinants of daily mobility affect male and female workers differently, finding persistent gender differences in travel patterns even when controlling for employment status and car access.
Salminen (2000) found that in Finland, men were involved in five out of every six traffic incidents during work hours, despite driving only three out of every four kilometers. The study also revealed that older workers (50-65 years) had the highest frequency of work-related traffic incidents, highlighting the importance of age as a factor in our analysis.
@@ 2.5 Work-Related incidents in Urban Contexts
Work-related traffic incidents, particularly those occurring during commutes, represent a significant subset of urban traffic incidents with unique characteristics and impacts. These incidents affect individual workers and have broader implications for urban mobility and economic productivity.
@@
Two years post-crash, 8% of previously unaffected car occupants in Sweden were on sickness absence or disability pension (Elrud et al., 2019). For cycle crashes, 20% of individuals had sickness absence, with duration varying by injury type and body region (Ohlin et al., 2018). These studies highlight the long-term impacts of traffic crashes on workers.
@@
Our study builds upon this body of work by explicitly examining work-related crashes in the context of Santiago, Chile, and evaluating the impact of a city-wide speed limit reduction on commuting safety. By focusing on this intersection of work-related mobility and urban traffic safety policies, and by analyzing the medical days of leave prescribed following crashes, collisions or falls, we contribute to a more nuanced understanding of how urban interventions can affect commuting safety and worker well-being in a Latin American context.
@@ 3 Context and Data
This study focuses on the Greater Santiago Metropolitan Area, the largest urban agglomeration in Chile. As of 2017, Santiago had a population of nearly 8 million inhabitants spread across an urban area of 867.75 square kilometers, comprising over 40 independent municipalities, with 45 within the urban core and the rest in the periphery. The city’s transportation system is characterized by an integrated multimodal network with a nearly flat fare structure across the metro, urban buses, and one rail service, allowing up to three transfers within a two-hour window (Muñoz et al., 2014).
The data for this study comes from Mutual de Seguridad de la Cámara Chilena de la Construcción (MUSEG), one of the institutions overseen by the Superintendencia de Seguridad Social (SUSESO) that prioritizes worker safety and insurance. MUSEG covers approximately 31% of all insured workers in Chile as of 2022. Our dataset spans a decade, encompassing 75.8K traffic crashes, collisions or falls in the greater Santiago Metropolitan Area from 2012 to 2021.
Each record contains:
@@
Fig. 1.a shows the monthly progression of event counts, indicating a positive trend, except for 2020, characterized by COVID-19 lockdowns.
The dataset contains non-fatal incidents only. We assess the severity of these incidents through the number of prescribed medical days of leave for each worker (see Fig. 1.b for its distribution, with a mean of 24.19 days in the whole dataset).
Figure 1: a) Monthly distribution of the number of crashes, collisions or falls. b) Histogram of the number of incidents concerning the number of medical leave days. c) Mean number of prescribed medical days of leave per month, with mean value annotations before and after Chile’s latest speed regulation law (August 4th, 2018).
A significant event within our study period is the implementation of a new speed limit regulation on August 4th, 2018, which reduced the urban speed limit from 60 km/h to 50 km/h (Gobierno de Chile, 2019). We hypothesize that this law impacted the severity of injuries from traffic crashes, collisions or falls. The data suggests so, as the mean number of medical days of leave decreased from 25.76 before the law was in effect to 22.44 after the law was in effect (see Fig. 1.c).
@@ 4.1 Data Representativeness
To understand traffic patterns in Santiago, we refer to the last official travel survey conducted in 2012, known as EOD2012 (SECTRA, 2015). This survey collected data on travel behavior, including trip origins, destinations, and modes of transportation, as well as demographic information of the resident population.
We compare origin-destination matrices built from both MUSEG and EOD2012 data. We calculate the Spearman correlation coefficient $r$ between the origin-destination pairs in EOD2012 and MUSEG data:
$r(X,Y)=\frac{\text{cov}(X_{R},Y_{R})}{\sigma_{X_{R}}\sigma_{Y_{R}}},$
cov is the covariance, and $\sigma$ is the standard deviation, applied to the rank-transformed versions of $X$ and $Y$, denoted by $X_{R}$ and $Y_{R}$. We estimate the correlation between OD matrices at the municipal level, as the EOD2012 data represents this level.
In particular, according to EOD2012, on an average working day in Santiago, there are 13.95 million trips, of which 18.7% are to-work trips, 2% are for-work trips, and 46% are to-home trips (with a fraction of these presumably mirroring work trips). We analyze the correlation between work trips.
@@ 4.2 Exploratory Analysis
We aggregate geolocated incidents and trip origins/destinations into a hexagonal grid with an edge length of approximately 1.4 km using the H3 system (Uber Technologies, Inc, 2024). We are interested in finding potential areas where mode-specific incidents are prevalent. As such, counting is insufficient, as some areas may have a small number of incidents, with a relatively high ratio of incidents in a given mode of transportation. Hence, we resorted to using a log-odds ratio with an uninformative Dirichlet prior (Monroe et al., 2008), a weighting schema that factors in the frequency and variability of incidents per cell. It is defined as follows:
$\frac{\hat{\delta}_{w}^{i-j}}{\sigma^{2}(\hat{\delta}_{w}^{i-j})},$
@@
$\hat{\delta}_{w}^{i-j}$ $=\log(\frac{y_{w}^{i}+\alpha_{w}}{n^{i}+\alpha_{0}-(y_{w}^{i}+\alpha_{w})})-\log(\frac{y_{w}^{j}+\alpha_{w}}{n^{j}+\alpha_{0}-(y_{w}^{j}+\alpha_{w})})$
@@
where $y_{w}^{i}$ is the observed number of incidents in mode $w$ within cell $i$, $n^{i}$ is the total number of incidents in cell $i$; conversely, here $j$ is interpreted as all other cells from the grid. The value of $\alpha_{w}$ is the prior parameter for the Dirichlet distribution, and it can be small to be uninformative, such as 0.01; $\alpha_{0}$ represents the sum of all prior parameters for each mode $w$, $\sum_{w}\alpha_{w}$.
4.3 Speed Limit Law Effects on Crash, Collision, or Fall Injury Severity
The days of medical leave represent a highly skewed variable (see Fig. 1.b), making it unsuitable for typical statistical tests to compare means. Therefore, we use regression analysis, where the dependent variable is the number of prescribed medical days of leave, and the independent variables are the characteristics of the trip/event.
While count data is usually modeled with Poisson regression, our data violates the Poisson distribution’s assumption that the conditional mean equals the conditional variance. We use a Negative Binomial (NB) regression model (Cameron and Trivedi, 1986) to account for the over-dispersion in our highly skewed medical leave data. The NB model probability density function is defined as:
$f(y_{i}\mid\mathbf{x}_{i})=\frac{\Gamma(y_{i}+\alpha)}{y_{i}!\leavevmode\nobreak\ \Gamma(\alpha)}\left(\frac{\alpha}{\alpha+u_{i}}\right)^{\tau}\left(\frac{u_{i}}{\alpha+u_{i}}\right)^{y_{i}},$
where $y_{i}$ is the medical days of leave, $\mathbf{x}_{i}$ is the set of predictor variables, $\Gamma$ is the gamma function, and $\alpha$ is the dispersion parameter. The conditional mean and variance are given by:
$E(y_{i}\mid\mathbf{x}_{i})=u_{i}=\exp(\mathbf{x}_{i}^{T}\mathbf{\beta});\text{Var}(y_{i}\mid\mathbf{x}_{i})=u_{i}\left(1+\frac{1}{\alpha}u_{i}\right).$

### v2

# Reducing Urban Speed Limits Decreases Work-Related Traffic Injury Severity: Evidence from Santiago, Chile

@@ Abstract
Work-related transportation incidents significantly impact urban mobility and productivity. These incidents include traffic crashes, collisions between vehicles, and falls that occurred during commuting or work-related transportation (e.g., falling while getting off a bus during the morning commute or while riding a bicycle for work). This study analyzes a decade of work-related transportation incident data (2012–2021) in Santiago, Chile, using records from a major worker’s insurance company. Using negative binomial regression, we assess the impact of a 2018 urban speed limit reduction law on incident injury severity. We also explore broader temporal, spatial, and demographic patterns in these incidents in urban and rural areas.
The urban speed limit reduction is associated with a decrease of 4.26 days in prescribed medical leave for incidents in urban areas, suggesting that lower speed limits contribute to reduced injury severity. Our broader analysis reveals distinct incident patterns across different groups. Workers traveling by motorcycle and bicycle experience more severe injuries when involved in traffic incidents, with marginal effects of 26.94 and 13.06 additional days of medical leave, respectively, compared to motorized vehicles. Women workers tend to have less severe injuries, with an average of 7.57 fewer days of medical leave. Age is also a significant factor, with older workers experiencing more severe injuries — each additional year of age is associated with 0.57 more days of medical leave. Our results provide insights for urban planning, transportation policy, and workplace safety initiatives.
Work trips are a crucial aspect of daily urban mobility, shaping the rhythm of cities and defining peak travel periods. The safety of these trips is important, as work-related traffic incidents (including crashes between vehicles, collisions with other road users, and falls during commuting or work transportation) impact individual health, urban productivity, and sustainability (United Nations, 2023). These incidents significantly impact urban dynamics by affecting workers, disrupting traffic, straining emergency services, and causing economic losses through reduced productivity (Chantith et al., 2021). In this paper, we use the term ‘incident’ to refer to traffic incidents that may result in injury or property damage, avoiding the term ‘accident’ which can imply unavoidability.
While disruptive to urban systems, traffic incidents are often underrepresented in transportation analyses. Traditional travel surveys rarely capture traffic incident data, and new data sources like mobile phone records, although helpful in understanding general mobility patterns (Widhalm et al., 2015; Graells-Garrido et al., 2021, 2023a), cannot identify specific traffic incidents due to anonymization and lack of meta-data. However, work-related incidents offer study opportunities, as they are more likely to be centrally recorded due to legal and insurance requirements.
@@
Various institutions prioritize worker safety and insurance in Santiago, overseen by the Superintendencia de Seguridad Social (SUSESO). This paper analyzes a dataset from one of these entities, Mutual de Seguridad de la Cámara Chilena de la Construcción (MUSEG), which covers approximately 31% of all insured workers in the country as of 2022. We quantify incident severity through prescribed medical days of leave — the time a medical professional determines a worker needs to recover before returning to work. This measure directly captures the immediate impact on workers’ lives and productivity. Our study examines 75.8K traffic incidents in the Santiago metropolitan area and surrounding rural regions (of which 33.4K could be geolocated with exact coordinates) from 2012 to 2021, revealing temporal, geographical, and demographic patterns.
A significant event within our study period is the implementation of a new speed limit regulation on August 4th, 2018, which reduced the urban speed limit from 60 km/h to 50 km/h (Gobierno de Chile, 2019). While this law directly regulates motorized vehicle speeds, its safety benefits extend to all road users, as most crashes involving pedestrians, cyclists, and motorcyclists also involve cars or other motor vehicles on shared roads. We use this policy change as a natural experiment, employing a Negative Binomial regression to assess its effect on incident severity.
@@
    What are the temporal, spatial, and demographic patterns of work-related transportation incidents in Santiago?
@@
    How did the 2018 speed limit reduction impact the severity of work-related transportation incidents?
@@ 1 Related Work
Research on urban transportation safety has shown that speed limit reductions can effectively reduce traffic injuries (Elvik, 2012; Cleland et al., 2020; Milton et al., 2021). However, less than 10% of road safety research focuses on low- and middle-income countries, despite these areas accounting for most road traffic deaths and injuries (Haghani et al., 2022). Our study addresses this gap by examining how speed limit changes affect work-related transportation incidents in a Latin American context, while also investigating the broader relationship between urban mobility patterns and traffic safety.
@@ 1.2 Spatial Analysis of Traffic incidents
In the Latin American context, Mangones et al. (2024) compared crash risk on arterial roads, Bus Rapid Transit (BRT), and Buses with High Level of Service (BHLS) corridors in Bogotá, Colombia. Their findings suggest that different public transport infrastructures have varying impacts on safety. BRT networks provide lower crash rates for less severe collisions but increase injuries and fatalities. This underscores the importance of considering different road types and transport modes in safety analyses. Sukhai et al. (2021), studying South Africa, another Global South country, found that contextual effects create unequal road safety risks, with higher vulnerabilities for pedestrians, women, and public transport users linked to rurality and socio-economic factors.
@@ 1.3 Impact of Speed Limit Laws
Research on speed limit reductions, particularly from the UK, provides strong evidence for their effectiveness. In Edinburgh, the introduction of 20 mph (32 km/h) limits resulted in a 22% reduction in collisions and 20% reduction in casualties compared to control zones (Kokka et al., 2024). A comprehensive meta-narrative synthesis found that 20 mph zones were consistently associated with reductions in collisions and casualties (Cleland et al., 2020). Multiple studies from Bristol and Edinburgh demonstrated that speed limit reductions led to measurable decreases in average traffic speeds and crashes (Pilkington et al., 2018; Nightingale et al., 2020).
While previous studies have shown the effectiveness of speed limit reductions in various contexts (Elvik, 2012; Milton et al., 2021; Cleland et al., 2020), including rigorous evaluations of citywide implementations (Kokka et al., 2024; Pilkington et al., 2018), research examining speed enforcement mechanisms provides additional supporting evidence. Montella et al. (2015) found that point-to-point speed enforcement on urban motorways reduced excessive speeding behaviors by over 80% and crashes by 32%, with even larger reductions during adverse conditions. Similarly, Goldenbeld and van Schagen (2005) documented a 21% reduction in both injury accidents and serious casualties after implementing targeted speed enforcement.
@@ 1.4 Gender and Age Dynamics in Traffic Incidents
González-Sánchez et al. (2021) investigated the influence of road type on traffic injury risk, considering gender and age. They found that men consistently have a higher risk of severe and fatal injuries across all modes of transport and road types. Havet et al. (2021) explored how determinants of daily mobility affect men and women workers differently, finding persistent gender differences in travel patterns even when controlling for employment status and car access.
Salminen (2000) found that in Finland, men were involved in five out of every six traffic incidents during work hours, despite driving only three out of every four kilometers. The study also revealed that older workers (50-65 years) had the highest frequency of work-related transportation incidents, highlighting the importance of age as a factor in our analysis.
@@ 1.5 Work-Related Incidents in Urban Contexts
Work-related transportation incidents, particularly those occurring during commutes, represent a significant subset of urban traffic incidents with unique characteristics and impacts. These incidents affect individual workers and have broader implications for urban mobility and economic productivity.
@@
Two years post-crash, 8% of car occupants in Sweden had permanent medical impairment (Elrud et al., 2019). For cycle crashes, 20% of individuals had sickness absence, with duration varying by injury type and body region (Ohlin et al., 2018). These studies highlight the long-term impacts of traffic crashes on workers.
@@
Our study builds upon this body of work by explicitly examining work-related crashes in the context of Santiago, Chile, and evaluating the impact of a city-wide speed limit reduction on commuting safety. By focusing on this intersection of work-related mobility and urban traffic safety policies, and by analyzing the prescribed medical days of leave following crashes, collisions or falls, we contribute to a more nuanced understanding of how urban interventions can affect commuting safety and worker well-being in a Latin American context.
@@ 2 Context and Data
This study focuses on the Greater Santiago Metropolitan Area, the largest urban agglomeration in Chile. As of 2017, Santiago had a population of nearly 8 million inhabitants spread across an urban area of 867.75 square kilometers, comprising over 40 independent municipalities. The city’s transportation system is characterized by an integrated multimodal network with a nearly flat fare structure across the metro, urban buses, and one rail service, allowing up to three transfers within a two-hour window (Muñoz et al., 2014).
The data for this study comes from Mutual de Seguridad de la Cámara Chilena de la Construcción (MUSEG), one of the institutions overseen by the Superintendencia de Seguridad Social (SUSESO) that manages worker’s insurance. Under Chilean law, this mandatory insurance covers any traffic incident (crash, collision, or fall) that occurs during a worker’s commute between home and workplace, between multiple workplaces, or during work-related transportation.
While police records are commonly used in traffic safety research, our insurance-based dataset offers unique advantages: 1) mandatory reporting of all non-fatal work incidents requiring medical attention, regardless of severity; 2) detailed medical outcomes through prescribed days of leave, providing a consistent measure of incident severity that captures impacts on workers’ lives; and 3) consistent recording over a decade. However, the dataset has several limitations: it excludes fatal incidents (handled through a different system), covers only 31% of insured workers with potential biases toward formal employment sectors, and like most incident databases, lacks exposure data (such as kilometers traveled) that would enable more detailed risk analysis. To address some of these limitations, we compared incident patterns with Santiago’s travel survey and excluded incidents with extremely long medical leave periods ($>$2 years) to avoid outlier effects (see Section 3 for details on data representativeness analysis).
Our dataset spans a decade, encompassing 75.8K work-related transportation incidents in the greater Santiago Metropolitan Area from 2012 to 2021, all occurring on public roads during either commuting trips (to/from work) or work-related travel. The geographical information in our records is organized by municipalities, which are the smallest administrative subdivisions in Chile (equivalent to districts or boroughs in other countries). Each record contains:
@@
Figure 1: a) Monthly distribution of the number of crashes, collisions or falls. b) Histogram of the number of incidents concerning the number of medical days of leave. c) Mean number of prescribed medical days of leave per month, with mean value annotations before and after Chile’s latest speed regulation law (August 4th, 2018).
Our dataset includes three types of incidents: 1) traffic crashes between vehicles; 2) collisions between vehicles and other road users; and 3) falls during transportation (e.g., while boarding/alighting public transport or cycling). All incidents in our analysis occurred on public roads during either commuting trips (to/from work) or work-related travel. Figure 1.a shows the monthly progression of event counts, indicating a positive trend, except for 2020, characterized by COVID-19 lockdowns.
The dataset contains non-fatal incidents only. We assess the severity of these incidents through the number of prescribed medical days of leave, reflecting the recovery time medical professionals deemed necessary for each injured worker (see Figure 1.b for its distribution, with a mean of 24.19 days in the whole dataset).
A significant event within our study period is the implementation of a new speed limit regulation on August 4th, 2018, which reduced the urban speed limit from 60 km/h to 50 km/h (Gobierno de Chile, 2019). We hypothesize that this law impacted the severity of injuries from traffic crashes, collisions or falls. The data suggests so, as the mean number of medical days of leave decreased from 25.76 before the law was in effect to 22.44 after the law was in effect (see Figure 1.c).
@@ 3 Methodology
Our methodology combines statistical and spatial analysis to assess the impact of Santiago’s 2018 speed limit reduction on work-related traffic incidents. We validate our dataset against the city’s 2012 travel survey, analyze incident patterns, and use negative binomial regression to quantify the speed limit law’s effect while controlling for transportation mode, timing, and demographics.
@@ 3.1 Data Representativeness
To understand whether the work trips that resulted in transportation incidents follow similar spatial patterns as regular work trips in Santiago, we refer to the last official travel survey conducted in 2012, known as EOD2012 (SECTRA, 2015). This survey collected data on travel behavior, including trip origins, destinations, and modes of transportation, as well as demographic information of the resident population. Using the worker’s municipality of residence as origin and municipality of workplace as destination, we calculate the Spearman correlation coefficient between origin-destination matrices built from both MUSEG and EOD2012 data:
$r(X,Y)=\frac{\text{cov}(X_{R},Y_{R})}{\sigma_{X_{R}}\sigma_{Y_{R}}},$ (1)
where cov is the covariance, and $\sigma$ is the standard deviation, applied to the rank-transformed versions of $X$ and $Y$, denoted by $X_{R}$ and $Y_{R}$. We estimate the correlation between OD matrices at the municipal level, as the EOD2012 data represents this level.
@@ 3.2 Exploratory Analysis
We aggregate geolocated incidents and trip origins/destinations into a hexagonal grid with an edge length of approximately 1.4 km using the H3 system (Uber Technologies, Inc, 2024). To measure spatial associations between areas and modes of transportation in incidents, we use a log-odds ratio with an uninformative Dirichlet prior (Monroe et al., 2008). The log-odds ratio identifies areas where specific modes of transportation are over- or under-represented in incidents. Areas with few incidents but a high proportion of a specific mode (e.g., 4 out of 5 incidents being cycling-related) could suggest meaningful patterns, but raw proportions and log-odds alone do not account for sampling variability. Using an uninformative Dirichlet prior provides the needed numerical stability in areas with few or no incidents of a particular mode, while adding minimal prior information to let patterns emerge naturally from the observations. The weighted log-odds ratio is defined as follows:
$\frac{\hat{\delta}_{w}^{i-j}}{\sigma^{2}(\hat{\delta}_{w}^{i-j})},$ (2)
@@
$\hat{\delta}_{w}^{i-j}$ $=\log(\frac{y_{w}^{i}+\alpha_{w}}{n^{i}+\alpha_{0}-(y_{w}^{i}+\alpha_{w})})-\log(\frac{y_{w}^{j}+\alpha_{w}}{n^{j}+\alpha_{0}-(y_{w}^{j}+\alpha_{w})})$ (3)
@@
where $y_{w}^{i}$ is the observed number of incidents in mode $w$ within cell $i$, $n^{i}$ is the total number of incidents in cell $i$; conversely, here $j$ is interpreted as all other cells from the grid. The value of $\alpha_{w}$, the prior parameter for the Dirichlet distribution, is set to 1 to be uninformative; $\alpha_{0}$ represents the sum of all prior parameters for each mode $w$, $\sum_{w}\alpha_{w}$.
3.3 Speed Limit Law Effects on Prescribed Medical Leave
Since all analyzed incidents occurred on public roads during commuting or work-related transportation, they were potentially affected by the speed limit change through either direct (vehicle speed) or indirect (general traffic calming) effects.
We use prescribed medical days of leave as a proxy measure for injury severity — more severe injuries typically require longer recovery periods and thus more medical leave. While other severity measures exist (like Abbreviated Injury Scale or hospitalization length), our dataset provides medical leave as the primary indicator of how severely an incident impacted a worker’s health and ability to return to work.
The days of medical leave represent a highly skewed variable (see Figure 1.b), making it unsuitable for typical statistical tests to compare means. Therefore, we use regression analysis, where the dependent variable is the number of prescribed medical days of leave, and the independent variables are the characteristics of the trip/event.
While count data is usually modeled with Poisson regression, our data violates the Poisson distribution’s assumption that the conditional mean equals the conditional variance. We use a Negative Binomial (NB) regression model (Cameron and Trivedi, 1986) to account for the over-dispersion in our highly skewed medical leave data. The NB model discrete probability density function is defined as:
$f(y_{i}\mid\mathbf{x}_{i})=\frac{\Gamma(y_{i}+\alpha)}{y_{i}!\leavevmode\nobreak\ \Gamma(\alpha)}\left(\frac{\alpha}{\alpha+u_{i}}\right)^{\alpha}\left(\frac{u_{i}}{\alpha+u_{i}}\right)^{y_{i}},$ (4)
where $y_{i}$ is the medical days of leave, $\mathbf{x}_{i}$ is the set of predictor variables, $\Gamma(\cdot)$ is the gamma function, and $\alpha$ is the dispersion parameter. The conditional mean and variance are given by:
$E(y_{i}\mid\mathbf{x}_{i})=u_{i}=\exp(\mathbf{x}_{i}^{T}\mathbf{\beta});\text{Var}(y_{i}\mid\mathbf{x}_{i})=u_{i}\left(1+\frac{1}{\alpha}u_{i}\right).$ (5)